Susan is five years older than Matt. In three years, she will be twice Matt's age.
How old are Susan and Matt?

Answer by sofiyac(983)

Matt=x years old
Susan=x+5
we need to multiply matt's age by 2 after 3 years in order to make their ages equal so we could solve for x
2(x+3)=(x+3+5)
2x+6=x+8
x=2
so matt is 2
and susan is 7